# A two-phase model of galaxy formation: III. The formation of globular clusters

### v1

@@ 4.4 Connection to host galaxies and halos
Figure 11: Total mass and number of GCs in galaxies at $z=0$. All results are shown as a function of host halo mass, $M_{\rm v}$. Top panel shows $M_{*}^{\rm(GCs)}$, the total mass of GCs within individual galaxies. Middle panel shows $\eta_{M}$, the ratio of $M_{*}^{\rm(GCs)}$ and $M_{\rm v}$. Bottom panel shows GC frequency, $T_{N}$, defined as the number of GCs, $N_{\rm GCs}$, per $10^{9}{\rm M_{\odot}}$ stellar mass of the host galaxy. Thick black curve is obtained by including all (in-situ + ex-situ) GCs within the central galaxy. Red and blue curves are obtained by only including Pop-I and Pop-II galactic GCs, respectively. Thin grey curve is obtained by only including in-situ galactic GCs. Grey dots show the results for galactic GCs of individual galaxies. Thick grey curve is obtained by including both galactic GCs and halo GCs (in satellite galaxies). For reference, green curves show $M_{*}$, the stellar mass of all stars (GC and non-GC), thin for in-situ and thick for in-situ plus ex-situ, both scaled by 5% for clarity. Shaded areas represent the 1-$\sigma$ ranges of the results. For comparison, purple dots show the observational results by Dornan & Harris (2023) for individual galaxies, and thick purple line shows their linear fit, $M_{*}^{\rm(GCs)}=3.0\times 10^{-5}M_{\rm v}$. Thin purple line shows the linear fit, $\log M_{\rm v}=4.15+\log M_{*}^{\rm(GCs)}$, obtained by Spitler & Forbes (2009). Purple squares show the results obtained by Peng et al. (2008) for individual galaxies from HST/ACS Virgo Cluster Survey, where we have converted their stellar mass to halo mass by the fitted relation obtained from our modeled galaxies. See §4.4 for a detailed discussion of this figure.
@@
The top panel of Fig. 11 shows $M_{*}^{\rm(GCs)}$, the mass of the GC population obtained by summing up the masses of individual GCs within the galaxy, as a function of halo mass, $M_{\rm v}$. Here, the analysis includes central galaxies at $z=0$ and the GCs survive in them. For convenience, the middle panel shows $\eta_{M}$, the GC mass fraction defined as the ratio between $M_{*}^{\rm(GCs)}$ and $M_{\rm v}$. For galactic GCs (in-situ + ex-situ; shown by the black curve), a linear relation, $M_{*}^{\rm(GCs)}\propto M_{\rm v}$, is seen over the entire range of halo mass considered here. The 1-$\sigma$ residual, as shown by the grey shaded area, is about $0.3\,{\rm dex}$, comparable to the root mean square error of $0.35\,{\rm dex}$ reported by Dornan & Harris (2023, see their §5). Including only the in-situ component decreases $M_{*}^{\rm(GCs)}$ by a halo-mass-dependent factor, from negligibly small at $M_{\rm v}\approx 10^{10}{\rm M_{\odot}}$, to about $0.5\,{\rm dex}$ at $10^{12}{\rm M_{\odot}}$ and about $1\,{\rm dex}$ at $\gtrsim 10^{14}{\rm M_{\odot}}$. Including additionally halo GCs (formed in satellite galaxies) increases $M_{*}^{\rm(GCs)}$ by a halo-mass-dependent factor, from negligibly small at $M_{\rm v}\lesssim 10^{12.5}{\rm M_{\odot}}$, to about $0.5\,{\rm dex}$ at $\gtrsim 10^{14}{\rm M_{\odot}}$, consistent with the results for the intra-cluster GCs obtained by Spitler & Forbes (2009) and Hudson et al. (2014). Separating galactic GCs into Pop-I and Pop-II components reveals the effects of metal enrichment on the two channels of GC formation, as already discussed in §4.2 using the mass function: below $M_{\rm v}\approx 10^{12.5}{\rm M_{\odot}}$, Pop-II GCs are dominating, while Pop-I GCs become comparable in abundance to and then slightly more abundant than Pop-II GCs above this halo mass. The difference between the two populations in the $M_{*}^{\rm(GCs)}$-$M_{\rm v}$ relation found here is very similar to the result obtained by Harris et al. (2015, see their figure 7), and we will quantify this in more detail in §4.5.
@@
The measurement of the $M_{*}^{\rm(GCs)}$-$M_{\rm v}$ relation in observations has to deal with a number of systematic uncertainties in, e.g., the selection of GCs, sample incompleteness, etc. Collecting data published in the literature, we found that observational results generally agree with each other within about $0.3\,{\rm dex}$, and the internal uncertainties originated from errors of individual data points are at a similar level (see e.g. figure 8 of Dornan & Harris, 2023). We therefore take two representative results from Spitler & Forbes (2009) and Dornan & Harris (2023), and show them in the first two panels of Fig. 11 by the thin and thick purple lines, respectively. For comparison, we also show results for individual galaxies obtained by Dornan & Harris (2023) using purple dots. The model prediction, shown by the black curve, matches the observation data well over the full range of the halo mass.
@@
The origin of the linear relation between $M_{*}^{\rm(GCs)}$ and $M_{\rm v}$ can be understood by comparing different curves in Fig. 11 produced by our model. Firstly, we note that, as shown by the green curves in the first two panels, the total stellar mass, $M_{*}$, is a non-linear function of $M_{\rm v}$ (see also observations of, e.g. Yang et al., 2012b; Behroozi et al., 2013), with $M_{*}/M_{\rm v}$ varying by more than $1\,{\rm dex}$ over the range of halo mass considered here. Thus, the relation between $M_{*}^{\rm(GCs)}$ and $M_{\rm*}$ must be non-linear, as clearly shown by the GC frequency, defined as the number of GCs per $10^{9}{\rm M_{\odot}}$ of the host galaxy, versus $M_{\rm v}$ in the bottom panel for galactic GCs predicted by our model and obtained by Peng et al. (2008) from observational data. The frequency of Pop-II (metal-poor) GCs closely follows that of the total population, with only small deviations above the MW mass. Pop-I is very different in that it follows the frequency of the total population only at the high end of the halo mass. The proximity of the Pop-II GCs to the total population suggests that the key to understanding the linear relation is to understand such a relation for Pop-II GCs. At $M_{\rm v}\lesssim 10^{12}{\rm M_{\odot}}$ where the in-situ Pop-II component dominates, the condition for a halo to host a significant number of Pop-II GCs is given by Eq. (65) and shown by the blue shaded area in Fig. 7. The negative correlation between $M_{\rm v}$ and $z$, which defines the boundary of this condition, leads to a narrower time window for a more massive halo to host GCs. It turns out that the metal enrichment of the ISM, which is positively correlated with galaxy mass and halo mass (Fig. 3), plays an important role in blocking the Pop-II channel of GC formation in massive halos and in shaping the linear relation between $M_{*}^{\rm(GCs)}$ and $M_{\rm v}$. At $M_{\rm v}\gtrsim 10^{12}{\rm M_{\odot}}$ where mergers become more frequent, this proportionality is preserved by the central limit theorem, as Pop-II GCs in satellite galaxies follow this linear relation and so does their summation. Thus, the origin of the observed $M_{*}^{\rm(GCs)}$-$M_{\rm v}$ relation is a combination of the initial linear relation set up by the metal enrichment process and the central limit theorem that preserves it.
@@
A similar conclusion for the importance of the central limit theorem in shaping the $M_{*}^{\rm(GCs)}$-$M_{\rm v}$ relation was reached by Boylan-Kolchin (2017) using a simple additive model for the formation of Pop-II GCs. In their model, an initial linear relation is assumed to populate tiny progenitor halos with GCs at $z\approx 6$, and the growth of a GC is followed by mass acquisition along its merger tree. Alternative explanations have also been proposed, e.g. by assuming a larger gas fraction and an enhanced GC formation rate, for halos with lower mass (e.g. Choksi et al., 2018; Choksi & Gnedin, 2019; El-Badry et al., 2019; Chen & Gnedin, 2022, 2023, 2024). Future observations to measure the gas fraction, density and metallicity, the reconstruction of the formation histories of GCs in low-$z$ galaxies (see e.g. Fig. 9), and to directly identify GC-forming galaxies at high $z$ (e.g. Vanzella et al., 2022, 2023; Welch et al., 2023; Lin et al., 2023; Claeyssens et al., 2023; Adamo et al., 2024; Messa et al., 2024; Fujimoto et al., 2024; Mowla et al., 2024), will be critical to distinguish different scenarios.
@@ 4.5 The metallicity dichotomy
Figure 12: Metallicity distribution of GCs. Three panels show the results for GCs in central galaxies with different halo masses at $z=0$. All galactic GCs (in-situ and ex-situ) are included in the analysis. Black, red and blue solid curves show the distributions of all, Pop-I and Pop-II GCs, respectively, survived until $z=0$. Black dashed curve is obtained by including all GCs formed in the history. For comparison, grey curve shows the metallicity distribution of all stars (GC and non-GC), obtained by using all sub-clouds formed in the history, weighted by their stellar masses. Distributions are arbitrarily normalized for clarity, but the relative normalizations for all, Pop-I and Pop-II GCs at $z=0$ are fixed. See §4.5 for the detailed description of this figure.
@@
To demonstrate the relative contribution to the GC formation of the two channels, in Fig. 12 we show the metallicity distribution of galactic GCs in central galaxies with different halo masses at $z=0$. The distribution for all survived GCs (black solid curve) is clearly bimodal, which is a consequence of the two channels of GC formation. For low-mass halos ($M_{\rm v}\sim 10^{11}{\rm M_{\odot}}$), Pop-II GCs (blue curve) dominates the total number, while Pop-I GCs (red curve) are just starting to appear in the metal-rich end of the distribution. For MW-size halos ($M_{\rm v}\sim 10^{12}{\rm M_{\odot}}$), The Pop-I channel becomes important, and the two populations are nearly equal in number. For more massive halos ($M_{\rm v}\gtrsim 10^{13}{\rm M_{\odot}}$), the relative contribution from the two channels appears to be stabilized, which is an outcome of the slightly increased Pop-I GC fraction at birth (black dashed curve) due to more enriched ISM, as compared to MW-size halos, and the stronger effect of tidal disruption due to the longer time for mass loss. For all halos and for both Pop-I and Pop-II, the metallicity of GCs is significantly lower than that of the total stellar component (grey curve), which reflects the preference of GC formation in a high-density environment in the early Universe (see Eqs. 63 and 65). Such a bimodal distribution of GCs has been found by observations in nearly all galaxies, using either direct metallicity measurements from spectra (e.g. Fahrion et al., 2020b), or conversion from color indices (e.g. Harris, 2023).
@@
as a function of halo mass. Both in-situ and ex-situ components are included in the analysis, and the results are shown for central galaxies at $z=0$. Our fiducial model (left panel) predicts a monotonically increasing Pop-I fraction with increasing halo mass (black thick curve). The fraction is near zero at $10^{10}{\rm M_{\odot}}$; increases rapidly to about $0.5$ at $10^{12}{\rm M_{\odot}}$; becomes saturated and slowly increases towards higher halo mass. This is consistent with the inferences from the effects of metal enrichment on the mass functions (§4.2), on the masses of GC systems (§4.4), and on the metallicity distributions (Fig. 12). The reason for the dominance of Pop-II GCs at $M_{\rm v}<10^{11}{\rm M_{\odot}}$ can also be inferred from the GC formation criteria shown in Fig. 7, where the mass assembly histories of halos with $M_{{\rm v},z=0}\lesssim 10^{11}{\rm M_{\odot}}$ do not intersect with the Pop-I channel (the region covered only by red). Similar results have been obtained by, e.g. Choksi & Gnedin (2019, see their figure 7) and El-Badry et al. (2019, see their figure 9). In more detail, we separate galactic GCs into in-situ and ex-situ components, and show their Pop-I fractions separately by the thin and dotted black curves, respectively. These two components show similar monotonic behavior to the total population in the Pop-I fraction, but with different amplitude and different halo mass for saturation. This difference can be explained by the fact that progenitors of satellites in general have lower halo mass than those of centrals at a given redshift, and their metal enrichment is systematically delayed, as seen by comparing the three panels of Fig. 2.
@@
For comparison, we plot the observational results from Harris (2023) for galaxies with a wide range of halo mass, with the BCGs highlighted by individual purple squares. For galaxies at $M_{\rm v}\gtrsim 10^{12.5}{\rm M_{\odot}}$, the Pop-I fraction produced by the fiducial model matches well the observation. However, the observed Pop-I fraction begins to decline at higher halo mass and drops to about $0.4$ at $10^{14}{\rm M_{\odot}}$, while the model prediction continues to increase. This significant discrepancy can arise from a number of sources. The first possibility is the systematic error in the observation. As many of the BCGs are quite distant, individual GCs within them are point sources even viewed by HST, and have to be distinguished from individual stars by, e.g., the color-magnitude diagram with trade-offs between completeness and contamination. The crowded environment in the inner regions of BCGs can also hide their GCs, and thus statistical inferences have to rely on the identification of the turnover point of the GC luminosity function and assumptions on the completeness curve. These systematics combined may lead to an underestimate of the Pop-I fraction in BCGs. The second possibility is the difference in the definition of Pop-I GCs between our model and the observation. The two populations in each BCG are obtained by Harris (2023) using a bi-Gaussian fitting to the distribution of the stellar $[{\rm Fe/H}]$ converted from a color index. In our model, on the other hand, the populations are classified using a physical criterion on the gas-phase oxygen abundance of the sub-clouds that host GCs (see §3.5). This difference in definition is not likely to be the main reason for the discrepancy, however, as the Pop-I fraction is related to the normalization of the two components of a bimodal distribution, rather than the metallicity of individual GCs. A linear conversion between oxygen and iron abundances for the alpha enhancement, such as that adopted by, e.g. Chen & Gnedin (2024, see their appendix), does not change the Pop-I fraction. We have also tried to use a bimodal fitting, instead of our physical criterion, to classify the two populations, and found no significant difference in the result. The final possibility is that the observed drop of the Pop-I fraction is physical and reflects a deficiency of our model, as we discuss in the following.
@@
One simplification of our model, as discussed in §3.4 of Paper-I, is the assumption of spherically symmetric accretion of gas into the SGC. Galaxies in the real Universe, however, can be fed by filamentary inflows of cold streams. The thin, cold streams of pristine gas are expected to penetrate the IGM, with only little mixing with the enriched outflow (see e.g. an analytical estimation by Li et al., 2023). High resolution numerical simulations (e.g. Mandelker et al., 2018) suggest that dense clumps can form within such streams, leading to even less mixing with the enriched gas. Thus, stream-fed galaxies may have a higher fraction of Pop-II GCs originated from cold-stream inflows. To test the effects of such accretion on GC formation, we design a suite of variants to the fiducial model by lowering the metallicity of sub-clouds in stream-fed galaxies. Formally, we define a galaxy to be stream-fed when it is in the fast phase of halo accretion, and if

### v2

@@ 4.4 Connection to host galaxies and halos
Figure 12: Total mass and number of GCs in galaxies at $z=0$. All results are shown as a function of host halo mass, $M_{\rm v}$. Top panel shows $M_{*}^{\rm(GCs)}$, the total mass of GCs within individual galaxies. Middle panel shows $\eta_{M}$, the ratio of $M_{*}^{\rm(GCs)}$ and $M_{\rm v}$. Bottom panel shows GC frequency, $T_{N}$, defined as the number of GCs, $N_{\rm GCs}$, per $10^{9}{\rm M_{\odot}}$ stellar mass of the host galaxy. Thick black curve is obtained by including all (in-situ + ex-situ) GCs within the central galaxy. Red and blue curves are obtained by only including Pop-I and Pop-II galactic GCs, respectively. Thin grey curve is obtained by only including in-situ galactic GCs. Grey dots show the results for galactic GCs of individual galaxies. Thick grey curve is obtained by including both galactic GCs and halo GCs (in satellite galaxies). For reference, green curves show $M_{*}$, the stellar mass of all stars (GC and non-GC), thin for in-situ only ($f_{\rm merge}=0$) and thick for in-situ plus ex-situ ($f_{\rm merge}=1$), both scaled by 5% for clarity. Shaded areas represent the 1-$\sigma$ ranges of the results. For comparison, purple dots show the observational results by Dornan & Harris (2023) for individual galaxies, and thick purple line shows their linear fit, $M_{*}^{\rm(GCs)}=3.0\times 10^{-5}M_{\rm v}$. Thin purple line shows the linear fit, $\log M_{\rm v}=4.15+\log M_{*}^{\rm(GCs)}$, obtained by Spitler & Forbes (2009). Purple squares show the results obtained by Peng et al. (2008) for individual galaxies from HST/ACS Virgo Cluster Survey, where we have converted their stellar mass to halo mass by the fitted relation obtained from our modeled galaxies. See §4.4 for a detailed discussion of this figure.
@@
The top panel of Fig. 12 shows $M_{*}^{\rm(GCs)}$, the mass of the GC population obtained by summing up the masses of individual GCs within the galaxy, as a function of halo mass, $M_{\rm v}$. Here, the analysis includes central galaxies at $z=0$ and the GCs survive in them. For convenience, the middle panel shows $\eta_{M}$, the GC mass fraction defined as the ratio between $M_{*}^{\rm(GCs)}$ and $M_{\rm v}$. For galactic GCs (in-situ + ex-situ; shown by the black curve), a linear relation, $M_{*}^{\rm(GCs)}\propto M_{\rm v}$, is seen over the entire range of halo mass considered here. The 1-$\sigma$ residual, as shown by the grey shaded area, is about $0.3\,{\rm dex}$, comparable to the root mean square error of $0.35\,{\rm dex}$ reported by Dornan & Harris (2023, see their §5). Including only the in-situ component decreases $M_{*}^{\rm(GCs)}$ by a halo-mass-dependent factor, from negligibly small at $M_{\rm v}\approx 10^{10}{\rm M_{\odot}}$, to about $0.5\,{\rm dex}$ at $10^{12}{\rm M_{\odot}}$ and about $1\,{\rm dex}$ at $\gtrsim 10^{14}{\rm M_{\odot}}$. Including additionally halo GCs (formed in satellite galaxies) increases $M_{*}^{\rm(GCs)}$ by a halo-mass-dependent factor, from negligibly small at $M_{\rm v}\lesssim 10^{12.5}{\rm M_{\odot}}$, to about $0.5\,{\rm dex}$ at $\gtrsim 10^{14}{\rm M_{\odot}}$, consistent with the results for the intracluster GCs obtained by Spitler & Forbes (2009) and Hudson et al. (2014). Separating galactic GCs into Pop-I and Pop-II components reveals the effects of metal enrichment on the two channels of GC formation, as already discussed in §4.2 using the mass function: below $M_{\rm v}\approx 10^{12.5}{\rm M_{\odot}}$, Pop-II GCs are dominating, while Pop-I GCs become comparable in abundance to and then slightly more abundant than Pop-II GCs above this halo mass. The difference between the two populations in the $M_{*}^{\rm(GCs)}$-$M_{\rm v}$ relation found here is very similar to the result obtained by Harris et al. (2015, see their figure 7), and we will quantify this in more detail in §4.5.
@@
The measurement of the $M_{*}^{\rm(GCs)}$-$M_{\rm v}$ relation in observations has to deal with a number of systematic uncertainties in, e.g., the selection of GCs, sample incompleteness, etc. Collecting data published in the literature, we found that observational results generally agree with each other within about $0.3\,{\rm dex}$, and the internal uncertainties originated from errors of individual data points are at a similar level (see e.g. figure 8 of Dornan & Harris, 2023). We therefore take two representative results from Spitler & Forbes (2009) and Dornan & Harris (2023), and show them in the first two panels of Fig. 12 by the thin and thick purple lines, respectively. For comparison, we also show results for individual galaxies obtained by Dornan & Harris (2023) using purple dots. The result produced by the model, shown by the black curve, matches the observational data well over the full range of the halo mass. This is partly because of our calibration with observations (see Appendix B), and partly because of the two underlying mechanisms in the model that shape the linearity between $M_{*}^{\rm(GCs)}$ and $M_{\rm v}$, as detailed below.
@@
The origin of the linear relation between $M_{*}^{\rm(GCs)}$ and $M_{\rm v}$ can be understood by comparing different curves in Fig. 12 produced by our model. Firstly, we note that, as shown by the green curves in the first two panels, the total stellar mass, $M_{*}$, is a non-linear function of $M_{\rm v}$ (see also observations of, e.g. Yang et al., 2012b; Behroozi et al., 2013), with $M_{*}/M_{\rm v}$ varying by more than $1\,{\rm dex}$ over the range of halo mass considered here. Thus, the relation between $M_{*}^{\rm(GCs)}$ and $M_{\rm*}$ must be non-linear, as clearly shown by the GC frequency, defined as the number of GCs per $10^{9}{\rm M_{\odot}}$ of the host galaxy, versus $M_{\rm v}$ in the bottom panel for galactic GCs produced by our model and obtained by Peng et al. (2008) from observational data. The frequency of Pop-II (metal-poor) GCs closely follows that of the total population, with only small deviations above the MW mass. Pop-I is very different in that it follows the frequency of the total population only at the high end of the halo mass. The proximity of the Pop-II GCs to the total population suggests that the key to understanding the linear relation is to understand such a relation for Pop-II GCs. At $M_{\rm v}\lesssim 10^{12}{\rm M_{\odot}}$ where the in-situ Pop-II component dominates, the condition for a halo to host a significant number of Pop-II GCs is given by Eq. (65) and shown by the blue shaded area in Fig. 8. The negative correlation between $M_{\rm v}$ and $z$, which defines the boundary of this condition, leads to a narrower time window for a more massive halo to host GCs. It turns out that the metal enrichment of the ISM, which is positively correlated with galaxy mass and halo mass (Fig. 3), plays an important role in blocking the Pop-II channel of GC formation in massive halos and in shaping the linear relation between $M_{*}^{\rm(GCs)}$ and $M_{\rm v}$. At $M_{\rm v}\gtrsim 10^{12}{\rm M_{\odot}}$ where mergers become more frequent, this proportionality is preserved by the central limit theorem, as Pop-II GCs in satellite galaxies follow this linear relation and so does their summation. Thus, the origin of the observed $M_{*}^{\rm(GCs)}$-$M_{\rm v}$ relation is a combination of the initial linear relation set up by the metal enrichment process and the central limit theorem that preserves it.
@@
A similar conclusion for the importance of the central limit theorem in shaping the $M_{*}^{\rm(GCs)}$-$M_{\rm v}$ relation was reached by Boylan-Kolchin (2017, see also ) using a simple additive model for the formation of Pop-II GCs. In their model, an initial linear relation is assumed to populate tiny progenitor halos with GCs at $z\approx 6$, and the growth of a GC is followed by mass acquisition along its merger tree. Alternative explanations have also been proposed, e.g. by assuming a larger gas fraction and an enhanced GC formation rate, for halos with lower mass (e.g. Choksi et al., 2018; Choksi & Gnedin, 2019; El-Badry et al., 2019; Chen & Gnedin, 2022, 2023, 2024). Future observations to measure the gas fraction, density and metallicity, the reconstruction of the formation histories of GCs in low-$z$ galaxies (see e.g. Fig. 10), and to directly identify GC-forming galaxies at high $z$ (e.g. Vanzella et al., 2022, 2023; Welch et al., 2023; Lin et al., 2023; Claeyssens et al., 2023; Adamo et al., 2024; Messa et al., 2024; Fujimoto et al., 2024; Mowla et al., 2024), will be critical to distinguish different scenarios.
@@ 4.5 The metallicity dichotomy
Figure 13: Metallicity distribution of GCs. Three panels show the results for GCs in central galaxies with different halo masses at $z=0$. All galactic GCs (in-situ and ex-situ) are included in the analysis. Black solid curve, red shade and blue shade show the distributions of all, Pop-I and Pop-II GCs, respectively, survived until $z=0$. Black dashed curve is obtained by including all GCs formed in the history. For comparison, grey curve shows the metallicity distribution of all stars (GC and non-GC), obtained by using all sub-clouds formed in the history, weighted by their stellar masses. Distributions are arbitrarily normalized for clarity. See §4.5 for the detailed description of this figure.
@@
To demonstrate the relative contribution to the GC formation of the two channels, in Fig. 13 we show the metallicity distribution of galactic GCs in central galaxies with different halo masses at $z=0$. The distribution for all survived GCs (black solid curve) is clearly bimodal, which is a consequence of the two channels of GC formation. For low-mass halos ($M_{\rm v}\sim 10^{11}{\rm M_{\odot}}$), Pop-II GCs (blue curve) dominates the total number, while Pop-I GCs (red curve) are just starting to appear in the metal-rich end of the distribution. For MW-size halos ($M_{\rm v}\sim 10^{12}{\rm M_{\odot}}$), The Pop-I channel becomes important, and the two populations are nearly equal in number. For more massive halos ($M_{\rm v}\gtrsim 10^{13}{\rm M_{\odot}}$), the relative contribution from the two channels appears to be stabilized, which is an outcome of the slightly increased Pop-I GC fraction at birth (black dashed curve) due to more enriched ISM, as compared to MW-size halos, and the stronger effect of tidal disruption due to the longer time for mass loss. For all halos and for both Pop-I and Pop-II, the metallicity of GCs is significantly lower than that of the total stellar component (grey curve), which reflects the preference of GC formation in a high-density environment in the early Universe (see Eqs. 63 and 65). Such a bimodal distribution of GCs has been found by observations in nearly all galaxies, using either direct metallicity measurements from spectra (e.g. Fahrion et al., 2020a), or conversion from color indices (e.g. Harris, 2023).
@@
as a function of halo mass. Both in-situ and ex-situ components are included in the analysis, and the results are shown for central galaxies at $z=0$. Our model predicts a monotonically increasing Pop-I fraction with increasing halo mass, regardless of the parameter choices. The fiducial model (left panel, black thick curve) is calibrated by the observation of Harris (2023, purple curve) at $M_{\rm v}\lesssim 10^{12.5}{\rm M_{\odot}}$ (see Appendix B), and thus matches the observation in this mass range. The Pop-I fraction is near zero at $10^{10}{\rm M_{\odot}}$; increases rapidly to about $0.5$ at $10^{12}{\rm M_{\odot}}$; becomes saturated and slowly increases towards higher halo mass. This is consistent with the inferences from the effects of metal enrichment on the mass functions (§4.2), on the masses of GC systems (§4.4), and on the metallicity distributions (Fig. 13). The reason for the dominance of Pop-II GCs at $M_{\rm v}<10^{11}{\rm M_{\odot}}$ can also be inferred from the GC formation criteria shown in Fig. 8, where the mass assembly histories of halos with $M_{{\rm v},z=0}\lesssim 10^{11}{\rm M_{\odot}}$ do not intersect with the Pop-I channel (the region covered only by red). Similar results have been obtained by, e.g. Choksi & Gnedin (2019, see their figure 7) and El-Badry et al. (2019, see their figure 9). In more detail, we separate galactic GCs into in-situ and ex-situ components, and show their Pop-I fractions separately by the thin and dotted black curves, respectively. These two components show similar monotonic behavior to the total population in the Pop-I fraction, but with different amplitude and different halo mass for saturation. This difference can be explained by the fact that progenitors of satellites in general have lower halo mass than those of centrals at a given redshift, and their metal enrichment is systematically delayed, as seen by comparing the three panels of Fig. 2.
@@
The observational results from Harris (2023) cover a wide range of halo mass, with the BCGs highlighted by individual purple squares. At $M_{\rm v}\approx 10^{12.5}{\rm M_{\odot}}$, the observed Pop-I fraction begins to decline and drops to about $0.4$ at $10^{14}{\rm M_{\odot}}$, while the result of fiducial model continues to increase. This significant discrepancy can arise from a number of sources. The first possibility is the systematic error in the observation. As many of the BCGs are quite distant, individual GCs within them are point sources even viewed by HST, and have to be distinguished from individual stars by, e.g., the color-magnitude diagram with trade-offs between completeness and contamination. The crowded environment in the inner regions of BCGs can also hide their GCs, and thus statistical inferences have to rely on the identification of the turnover point of the GC luminosity function and assumptions on the completeness curve. These systematics combined may lead to an underestimate of the Pop-I fraction in BCGs. The second possibility is the difference in the definition of Pop-I GCs between our model and the observation. The two populations in each BCG are obtained by Harris (2023) using a bi-Gaussian fitting to the distribution of the stellar $[{\rm Fe/H}]$ converted from a color index. In our model, on the other hand, the populations are classified using a physical criterion on the gas-phase oxygen abundance of the sub-clouds that host GCs (see §3.5). This difference in definition is not likely to be the main reason for the discrepancy, however, as the Pop-I fraction is related to the normalization of the two components of a bimodal distribution, rather than the metallicity of individual GCs. A linear conversion between oxygen and iron abundances for the alpha enhancement, such as that adopted by, e.g. Chen & Gnedin (2024, see their appendix), does not change the Pop-I fraction. We have also tried to use a bimodal fitting, instead of our physical criterion, to classify the two populations, and found no significant difference in the result. The final possibility is that the observed drop of the Pop-I fraction is physical and reflects a deficiency of our model, as we discuss in the following.
@@
One simplification of our model, as discussed in §3.4 of Paper-I, is the assumption of spherically symmetric accretion of gas into the SGC. Galaxies in the real Universe, however, can be fed by filamentary inflows of cold streams. The thin, cold streams of pristine gas are expected to penetrate the IGM, with only little mixing with the enriched outflow (see e.g. an analytical estimation by Li et al., 2024). High resolution numerical simulations (e.g. Mandelker et al., 2018) suggest that dense clumps can form within such streams, leading to even less mixing with the enriched gas. Thus, stream-fed galaxies may have a higher fraction of Pop-II GCs originated from cold-stream inflows. To test the effects of such accretion on GC formation, we design a suite of variants to the fiducial model by lowering the metallicity of sub-clouds in stream-fed galaxies. Formally, we define a galaxy to be stream-fed when it is in the fast phase of halo accretion, and if